OpenStudy (anonymous):

Describe the vertical asymptote(s) and hole(s) for the graph of y= (x+3)(x+4) / (x+4)(x+2). Answers: a) asymptote: x=3 and hole: x=-2 b) asymptote: x=-2 and hole: x= -4 c) asymptote: x=-2 and hole: x=4 d) asymptote: x=2 and hole: x=4

OpenStudy (anonymous):

@nincompoop @Hero please help me

OpenStudy (hayleymeyer):

cross off what you Dont think is the answer

OpenStudy (anonymous):

I WILL FAN + MEDAL PLEASE HELP !

OpenStudy (anonymous):

@dumbcow @hick4life

Nnesha (nnesha):

for \(\color{green}{\rm Vertical~ asy.}\) set the denominator equal to zero and then solve for the variable. for\(\color{green}{\rm Horizontal ~asy.}\) focus on highest degrees ~if the highest degree of the numerator is greater than the denominator then `No horizontal asy.` \[\color{reD}{\rm N}>\color{blue}{\rm D}\] example \[\large\rm \frac{ 7x^\color{ReD}{3} +1}{ 4x^\color{blue}{2}+3 }\] ~if the highest degree of the denominator is greater than the highest degree of the numerator then `y=0` would be horizontal asy. \[\rm \color{reD}{N}<\color{blue}{\rm D}\] example:\[\large\rm \frac{ 7x^\color{red}{2}+1 }{ 4x^\color{blue}{3}+3 }\] ~if both degrees are the same then divide the leading coefficient of the numerator by the leading coefficient of the denominator \[\rm \color{red}{N}=\color{blue}{D}\] \[\large\rm \frac{ 8x^\color{reD}{3}+1 }{ 4x^\color{blue}{3}+3 }\] \[\rm \frac{ 8x^3 }{ 4x^3 } =2\] horizontal asy. =2

OpenStudy (hick4life):

ok lets see what we got here

OpenStudy (anonymous):

i will medal and fan if you get me the answer @hick4life

Nnesha (nnesha):

\[\rm f(x)=\frac{ (x+3)(x+4)}{ (x+2)(x+4)}\] first simplify is there anything you can cacnel out?

Nnesha (nnesha):

cancel*

OpenStudy (hick4life):

Do we have a graph line

OpenStudy (dumbcow):

@jordanaz26 , just getting the answer wont help you on the next question...

OpenStudy (anonymous):

yeah we can cancel out (x+4) @Nnesha

Nnesha (nnesha):

yes right! so \[f(x)=\frac{ (x+3) }{ x+2}\]set the denominator equal to zero to find vertical asy

Nnesha (nnesha):

x+2=0 solve for x

OpenStudy (hick4life):

Go ahead and cancel (4x+)

OpenStudy (anonymous):

x=-2 @Nnesha

Nnesha (nnesha):

:=) looks good wait a sec plz

OpenStudy (anonymous):

sure ! so we got the vertical asymptote which is x:-2, but it also asks for the hole(s) @Nnesha

OpenStudy (hick4life):

Its going to be B.

Nnesha (nnesha):

yeah i'mm trying to find hole in my notebook i forgot how to find it

Nnesha (nnesha):

ahh okay what are the factors that in common to both numerator and denominator ?

OpenStudy (anonymous):

of 3 and 2 ? @Nnesha

Nnesha (nnesha):

common factors would be hole \[\rm f(x)=\frac{ (x+3)\color{reD}{(x+4)}}{ (x+2)\color{red}{(x+4)}}\] in this question x+4 is common right so set it equal to zero x+4=0 solve for x that value would be hole

OpenStudy (anonymous):

oh !!! so its x=-4 ! thank you so much ! fan and medal ! @Nnesha

Nnesha (nnesha):

my pleasure

Nnesha (nnesha):

and yes it's -4